OpenStudy (anonymous):

Square root of 4z + 1 equals 3 plus the square root of 4z - 2. How to do?

OpenStudy (anonymous):

\[\sqrt{4z+1} = 3 + \sqrt{4z-2}\]

OpenStudy (anonymous):

No solutions. You square both sides to get rid of the square root on the 4z+1. Simplify. You get 4z+1 = 9+6*sqroot(4x+2) + 4z -2 Subtract 4z from both sides. 1 = 9 + 6* sqroot (4z+2) - 2 1 = 7 + 6* sqroot (4z+2) -1 = sqroot (4z+2)